Skip to main contentSkip to solution

Two thirds of the students in a class are boys and the rest are girls. It is known that the probability of a girl getting a first class is 0.25 and that of a boy getting a first class is 0.28, then the probability that a student chosen at random will get first class marks in the subject is

Solution

✅ Correct Option: 2

By total probability, P=23(0.28)+13(0.25)=0.1867+0.0833=0.27P = \frac{2}{3}(0.28) + \frac{1}{3}(0.25) = 0.1867 + 0.0833 = 0.27.

Keyboard Shortcuts

  • Left arrow: Previous question
  • Right arrow: Next question
  • S key: Jump to solution
  • Q key: Jump to question